### v3

# One Demo Is All It Takes: Planning Domain Derivation with LLMs from A Single Demonstration

Abstract
Pre-trained large language models (LLMs) show promise for robotic task planning but often struggle to guarantee correctness in long-horizon problems. Task and motion planning (TAMP) addresses this by grounding symbolic plans in low-level execution, yet it relies heavily on manually engineered planning domains. To improve long-horizon planning reliability and reduce human intervention, we present Planning Domain Derivation with LLMs (PDDLLM), a framework that automatically induces symbolic predicates and actions directly from demonstration trajectories by combining LLM reasoning with physical simulation roll-outs. Unlike prior domain-inference methods that rely on partially predefined or language descriptions of planning domains, PDDLLM constructs domains without manual domain initialization and automatically integrates them with motion planners to produce executable plans, enhancing long-horizon planning automation. Across 1,200 tasks in nine environments, PDDLLM outperforms six LLM-based planning baselines, achieving at least 20% higher success rates, reduced token costs, and successful deployment on multiple physical robot platforms.
1 Introduction
Robotic planning remains challenging in complex scenarios that require abstract, long-horizon reasoning. Large language models (LLMs) demonstrate strong generalization in this domain but often struggle with temporal dependencies in extended tasks Huang et al. (2022, 2023). Task and Motion Planning (TAMP) frameworks provide robustness in long-horizon reasoning by integrating high-level symbolic reasoning with low-level motion planning. However, they face two major limitations: (i) planning domains, expressed in symbolic planning languages such as PDDL (McDermott et al., 1998), are difficult to ground in complex geometric information, and (ii) these domains are labor-intensive to manually construct Garrett et al. (2020, 2021); Khodeir et al. (2023); Silver et al. (2020). Recent vision-language-action models have advanced the first challenge by improving geometric grounding through semantic action instructions, but they pay limited attention to long-horizon reasoning Black et al. (2024); Kim et al. (2024); Zitkovich (2023); Li et al. (2024); Team (2024). Complementarily, our work focuses on solving the second challenge by reducing the human effort required for domain construction, thereby enabling automated and scalable long-horizon reasoning.
Existing domain-generation methods typically rely on predesigned elements (predicates or actions) to complete a domain Silver et al. (2023); Kumar et al. (2023), where manual predesign remains critical and often depends on curated training data Huang et al. (2025). LLM-based approaches also demand extensive natural-language descriptions of PDDL domains and careful prompt engineering Guan et al. (2023a); Oswald et al. (2025). Meanwhile, recent advances in world-model learning show that LLMs can serve as compact explainers and reasoners for robot behavior and object relations (Guan et al., 2023b; Zhao et al., 2024; Liang et al., 2024; Tang et al., 2024). Motivated by this, we propose to combine LLM reasoning with physical simulation to generate planning domains, eliminating reliance on manual predesign, curated training data, and detailed textual descriptions.
We introduce PDDLLM, the first method to construct planning domains without manual predesign, training data, or extensive human descriptions. Moreover, PDDLLM automates integration with low-level motion planners, further reducing the need for manual intervention. Our method targets at long-horizon planning problems, aligning with the scope of TAMP studies. Managing action sequences and task structures is essential for these tasks such as household activities and puzzle solving. The main contributions of this paper are:
  • An algorithm that combines LLMs and physical simulation to automatically generate a human-interpretable planning domain from a single demonstration.
  • A Logical Constraint Adapter (LoCA) that systematically interfaces the generated domain with low-level motion planners.
  • An extensive evaluation on over 1,200 tasks in nine environments, demonstrating superior long-horizon planning performance and more efficient token usage than state-of-the-art baselines, with successful deployment on three physical robot platforms.
2 Related Work
Task planning with pre-trained large models
With the advent of pre-trained large models (LMs), the use of LLMs and vision language models (VLMs) has significantly advanced the performance of task planners (Huang et al., 2023; Wang et al., 2024a; Chen et al., 2024; Li et al., 2023). Although many studies have demonstrated the generalization capabilities of LM-based task planners, they often lack robustness and struggle with long-horizon tasks that require complex reasoning (Wang et al., 2024a; Sermanet et al., 2023; Driess et al., 2023; Brohan et al., 2023; Huang et al., 2022). To address this limitation, recent research has explored guiding task planners with LM-derived heuristics to accelerate informed search. These approaches integrate symbolic search with LMs to accelerate task planning and reduce search complexity. Notable efforts include heuristics for prioritizing feasible states (Zhao et al., 2024), ranking feasible actions (Yang et al., 2024; Meng et al., 2024; Hu et al., 2023; Zhao et al., 2024), and pruning search trees (Silver et al., 2024). However, a major limitation of these methods lies in their reliance on manually constructed symbolic planning domains to build search trees, which imposes additional development overhead and reduces flexibility.
Learning planning domains
A recent line of research aims to infer the planning domain directly from human demonstrations, environment interactions, or natural language. However, these approaches often depend on partially or fully predefined symbolic predicates and actions as well as extensive training dataset (Silver et al., 2023; Kumar et al., 2023; Huang et al., 2025; Wong et al., 2023; Liu et al., 2025). Some recent studies have explored leveraging LMs for domain generation, primarily by extending manually defined domains with additional predicates and actions (Liang et al., 2024; Athalye et al., 2024; Byrnes et al., 2024). The approaches proposed by Guan et al. (2023b); Han et al. (2024) generates planning domains requiring prompts containing human-crafted planning examples or detailed descriptions of predicates and actions. Moreover, many of these studies assume that symbolic actions must be manually linked to existing motion-level skills (Kumar et al., 2023; Huang et al., 2025; Liang et al., 2024; Athalye et al., 2024), requiring labor-intensive alignment between planning domains and low-level motion planners. LoCA addresses this limitation by automatically grounding symbolic actions into motion constraints, thereby reducing manual effort and enhancing scalability and autonomy.
3 Preliminaries
PDDL is a standard formal language used to specify planning problems. The object set $\mathcal{O}$ represents the environment’s objects, whose continuous state $\mathcal{S}$, such as pose, color, and size, can be queried via a perception function $\mathcal{I}:\mathcal{O}\times\mathcal{I}\rightarrow\mathcal{S}$. The PDDL domain $\mathcal{D}=(\mathcal{P},\mathcal{A})$ describes the general rules of the environment, consisting of a set of logical predicates $\mathcal{P}$ and a set of logical actions $\mathcal{A}$. A logical predicate $p\in\mathcal{P}$ specifies either intrinsic properties of an object $o$ or relations between objects (e.g.,(cooked ?o1), (on ?o1 ?o2)). Each predicate is evaluated by a binary classifier over the continuous state, returning true or false. A symbolic descripton of the environment $\mathcal{X}$ can be obainted by grounding $\mathcal{P}$ across $\mathcal{S}$ (i.e. $\mathcal{S}\times\mathcal{P}\rightarrow\mathcal{X}$). A logical action $a\in\mathcal{A}$ consists of a precondition $\mathcal{P}_{pre}$ $\langle p_{1},p_{2},\dots\rangle$ and an action’s effect $\mathcal{P}_{eff}=\langle p_{1}^{\prime},p_{2}^{\prime},\dots\rangle$. The precondition represents a set of predicates that must be satisfied for the action to be executed, while the action’s effect describes the change of the resultant state upon action completion. Logical actions define the logical state transitions $\mathcal{X}^{(t)}\times a^{(t)}\rightarrow\mathcal{X}^{(t+1)}$. Thus, any planning problem $\langle\mathcal{S}^{(init)},\mathcal{S}^{(final)}\rangle$ can be formulated as a logical planning problem $\mathcal{Q}=\langle\mathcal{O},\mathcal{D},\mathcal{X}^{(init)},\mathcal{X}^{(final)}\rangle$, which is solved by a symbolic planner to produce a task plan ${a^{(0)},a^{(1)},\dots,a^{(T-1)}}=\texttt{PDDLSolver}(\mathcal{Q})$. Each logical action $a$ must then be integrated with corresponding motion planning skills to generate continuous robotic actions $\tilde{a}$ for execution.
4 Problem Statement
Our system solves a new long-horizon planning problem given an expert demonstration trajectory $\tau_{demo}$ and its associated task description $T_{demo}$. The trajectory $\tau_{demo}$ is a sequence of continuous environment states $\mathcal{S}^{(0)},\mathcal{S}^{(1)},\dots$}, while $T_{demo}$ is a short natural language phrase describing the task. Following prior work (Silver et al., 2023; Kumar et al., 2023; Huang et al., 2025), the robot is the sole acting agent during data collection.
Given a new planning problem, PDDLLM derives a PDDL domain from $(T_{demo},\tau_{demo})$ which generates a sequence of symbolic actions as the task plan. LoCA, a component within PDDLLM, interfaces the task plan with an off-the-shelf motion planner, which produces executable robot trajectories. Formally, the problem can be expressed as:
$\tilde{a}^{(0)},\tilde{a}^{(1)},\dots,\tilde{a}^{(L-1)}$ $=MotionPlanner(PDDLLM(\mathcal{S}^{(init)}_{new},\mathcal{S}^{(goal)}_{new},T_{demo},\tau_{demo}))$ (1)
where $\mathcal{S}_{new}^{(init)}$ and $\mathcal{S}_{new}^{(goal)}$ define the initial and goal states of the new problem, $L$ denotes the task plan length, and $\tilde{a}^{(0:L)}$ indicates an executable robot trajectory.
5 Methodology
Figure˜1 presents an overview of the PDDLLM framework. With $T_{demo}$ and $\tau_{demo}$ as inputs, PDDLLM constructs a relevant predicate library through predicate imagination and generates an action library via action invention. Ultimately, these predicate and action libraries are compiled into an executable PDDL planning domain, automatically interfaced with motion planners via LoCA. In the following sections, we provide a detailed explanation of each step in the framework.
Figure 1: Overview of the proposed framework. (1) Human demonstrations, in the form of manipulation trajectories, and the corresponding task descriptions, serve as input. Implementation details is shown in Section˜A.12. (2) PDDLLM initiates thousands of parallel simulations to imagine predicates by summarizing the simulation roll-outs with LLMs, returning the predicate library along with the relevance of each predicate to the current task. (3) Actions are invented by an LLM that summarizes logical state transition patterns from the demonstration, which is grounded into logical states using the imagined predicates. (4) The predicates and actions are compiled into a planning domain, which is automatically interfaced with motion planning algorithm by the Logical Constrain Adapter (LoCA) to solve new tasks.
5.1 PDDL Domain Generation
Given a human demonstration and the corresponding task description, we combine simulated physical interaction and LLM reasoning to produce an executable PDDL planning domain through predicate imagination and action invention.
5.1.1 Predicate Imagination
Predicate imagination refers to the process of summarizing simulated object relation roll-outs into meaningful predicates. The process consists of two stages: Stage 1 generates first-order predicates and Stage 2 further derives the higher-order predicates from the first-order predicates.
Stage 1. In this stage, PDDLLM generates first-order predicates, which directly describe the physical properties or relations of the objects (e.g., (is_on ?o1 ?o2), (smaller ?o1 ?o2)), by summarizing simulated object interactions using an LLM.
Definition 1 (Feature Space).
The feature space is defined as a set of continuous state variables, such as position, orientation, size, and color, that fully characterize the state of each object in the environment.
Following Definition˜1, the feature space is defined as a set of variables, such as Cartesian coordinates $(x,y,z)$ for object positions and RGB values $(r,g,b)$ for colors. The feature space is bounded by real-world constraints. Object states are uniformly sampled across this space to span a diverse range of object-object and object-environment interactions. Each sampled state undergoes a two-step verification process. First, it is evaluated for physical feasibility using a physical simulator, which serves as a physical knowledge base to capture complex dynamics beyond the reasoning capacity of LLMs. Only physically valid states are retained.
Next, PDDLLM partitions the feature space into a finite collection of subspaces. The range of each feature is divided into intervals, with the length of each interval being a hyperparameter. The intersections of these intervals specify the subspaces. Each object state can be mapped into one of the subspaces. Each subspace is analyzed to determine whether it contains feasible object states, as verified through simulation in the previous verification step. If so, simulation roll-out summaries are generated as prompts following a predefined scene-description template. Prompt generation is automated as it only requires the replacement of some keywords (such as “position” with “color”) and the specification of interval boundaries. An LLM is then prompted to summarize subspaces into meaningful predicates and select those relevant to the task. The subspace boundaries serve as predicate physical constraints, enabling the classification of whether a predicate holds true. Figure˜2.a illustrates an example of predicate generation for positional relations between objects.
Figure 2: a. This example illustrates the imagination of predicates for relative object positions. Let $u$ be a configurable variable determined by the feature subspace size. Object poses are sampled and simulated, with infeasible cases filtered out. Feasible subspaces are provided to the LLM to generate first-order predicates with their corresponding physical constraints. Higher-order predicates can be further derived using logical operators (e.g., "not", "for all") from first-order predicates. Diverse predicate examples are provided in Section˜A.9 b. This example shows how the Stack action is invented. Continuous states are grounded into logical states using the imagined predicates, where the state transition represents the logical action. By prompting the LLM with the pair of the current state and the next state after the transition, we obtain the PDDL definition of the action Stack. c. The integration of actions with the motion planner is handled automatically by LoCA, which retrieves the physical constraints associated with each first-order predicate in the action effect set $\mathcal{P}_{eff}$ and applies these constraints for motion planning. Prompt examples for a and b are provided in Section˜A.3.
Stage 2. First-order predicates are combined with logical operators and quantifiers to construct higher-order predicates. They are systematically combined in all possible ways to generate more complex logical expressions. Following prior work (Curtis et al., 2022; Silver et al., 2023), we primarily use the negation operator, along with the quantifiers “for all” ($\forall$) and “there exists” ($\exists$). For example, (is_on ?o1, ?o2) can be negated to produce (not_is_on ?o1, ?o2). When combined with the universal quantifier, this further yields ($\mathrm{}{\forall}$_o1_not_is_on ?o1, ?o2), meaning that for any o1 in the environment, (is_on ?o1, ?o2) is false. This indicates that object o2 is on top.
5.1.2 Action Invention
After constructing the predicate library, the human demonstration $\mathbf{\tau}_{demo}$ can be mapped into the logical space as $\mathbf{\tau}_{demo}^{logic}$ by grounding all relevant predicates at each time step. The logical state transitions within $\mathbf{\tau}_{demo}^{logic}$ signify the execution of actions. An advantage of learning actions in the logical space is that it simplifies pattern recognition by focusing on moments of logical state transitions, effectively transforming long trajectories into concise logical representations. For instance, the continuous manipulation trajectory $\mathbf{\tau}_{demo}$ in Figure˜2.b contains over 1000 time steps while $\mathbf{\tau}_{demo}^{logic}$ is reduced to merely 2 steps. Logical state transitions are extracted from pre-and-post-change states and presented to the LLM as prompts. After experimenting with various prompt structures, we found this direct, structured method to be the most effective. A concrete example of action invention for Stack is elaborated in Figure˜2.b.
5.1.3 Generation Refinement
To improve the quality and accuracy of the generated planning domains, we adopt a parallel prompting strategy. Specifically, multiple candidate PDDL models are generated from the same demonstration, and candidates that fail to produce feasible task plans are replaced by alternatives. Empirically, we find that when the number of parallel prompts exceeds five, the correctness of the final domain converges to a stable level. Examples of the parallel prompting procedure are shown in Section˜A.7
5.2 Automatic Interfacing with Motion Planner
Combining the PDDL planning domain with symbolic solvers yields a task plan $a^{(0)},a^{(1)},\dots,a^{(T-1)}$}. The next step is to establish an interface between each logical action and the low-level motion planner.
In PDDLLM, this interfacing is fully automated through the Logical Constraint Adapter (LoCA). Prior approaches relied on manually encoding logical actions as mathematical constraints for motion planning (Toussaint, 2015). LoCA automates and generalizes this process by directly retrieving the physical constraints associated with the first-order predicates in an action’s effect set $\mathcal{P}_{\text{eff}}$. These constraints, expressed as mathematical inequalities, are sequentially applied to the motion planner, ensuring that the required predicates are satisfied once the action is executed. In this way, LoCA automatically transforms each logical action into a standard constrained motion planning problem. As a result, the generated trajectory is guaranteed to align with the semantic meaning of the action, as illustrated in Figure˜2.c. This eliminates the need for manually engineered interfaces or predefined motion-planning skills, thereby streamlining the integration between symbolic planning and low-level control.
Additionally, our framework is also compatible with recent advances such as vision–language–action (VLA) models by directly using logical actions as prompts. Experiments in Section˜8 demonstrate PDDLLM’s ability to automatically integrate with both traditional and learning-based motion planning.
6 Experiments and Baselines
Building on evaluation practices in TAMP, we consider tasks with concrete conditions such as object position, orientation, size, and color (Silver et al., 2023; Garrett et al., 2020; Kumar et al., 2023; Liang et al., 2024; Huang et al., 2025). We structure our experiments along three dimensions, Task Diversity, Task Complexity, and Knowledge Transferability, whose definitions are provided in Section˜A.10.1, with task details described in Section˜A.4. Our experiments are conducted in PyBullet Simulation (Coumans and Bai, 2021), with the symbolic solver from PDDLStream (Garrett et al., 2020) and the motion planning algorithm from PyBullet-Planning (Garrett et al., 2015; Garrett, 2018). The LLM we used is GPT-4o (OpenAI, 2023).
6.1 Baselines and Ablations
We implement six baselines and one ablation of PDDLLM to comprehensively evaluate our method. GPT-4o is used as the default LLM unless otherwise specified, and the motion planning algorithm is kept the same for all methods. The complete details are provided in Section˜A.2
  • LLMTAMP, o1-TAMP, R1-TAMP: LLM-based task and motion planning (LLMTAMP) inspired by Huang et al. (2022); Li et al. (2022), which use LLMs for task planning, with language description of planning task as input. The task execution demonstration, same as those used in PDDLLM, was provided in the form of natural language in prompt. In addition to GPT-4o, reasoning LLMs, OpenAI’s o1 (OpenAI, 2024) and Deepseek’s R1 (DeepSeek-AI, 2025), are used as backbones for o1-TAMP and R1-TAMP, respectively.
  • LLMTAMP-$F$: Following the method by Huang et al. (2023); Chen et al. (2023), LLMTAMP-$F$ extends LLMTAMP with a failure feedback loop.
  • LLMTAMP-FR: Following Wang et al. (2024b), LLMTAMP-FR extends failure detection by providing specific failure reasons to guide replanning with the LLM.
  • Expert Design: The expert design baseline uses expert-crafted planning domains with symbolic solvers. Expert-designed domains are refined from PDDLLM-derived domains by an expert.
In addition to the six baselines, we include an ablation of our method, RuleAsMem, which uses the generated PDDL domain as contextual memory for LLM task planners, replacing symbolic solvers.
Robot planning is required to be real-time in robot deployment, imposing constraints on the planning time allowance (Gammell et al., 2015; Garrett et al., 2015, 2020). In our experiment, a uniform planning time limit of 50 seconds is applied to all planning problems and methods. We measure performance using the planning success rate, as in other studies (Huang et al., 2025; Silver et al., 2023; Kumar et al., 2023). The planning time and token cost are used to measure the planning cost (Zhong et al., 2024). Three parallel runs were conducted to compute the mean and standard error for the planning success rate.
Table 1: Planning success rate (%) across tasks for all methods (time limit = 50 s). The best results are highlighted in bold. Expert is excluded from the comparison, as it requires additional manual effort and serves as an upper bound.
Method Expert LLMTAMP LLMTAMP-$F$ LLMTAMP-FR RuleAsMem PDDLLM
Stack $98.5\pm 0.8$ $41.7\pm 4.3$ $70.8\pm 1.4$ $64.2\pm 3.1$ $85.5\pm 2.9$ 97.5 $\pm$ 1.6
Unstack $100\pm 0.0$ $89.4\pm 1.5$ $94.6\pm 0.9$ $92.1\pm 2.3$ $88.4\pm 1.2$ 97.7 $\pm$ 0.7
Color Classification $100\pm 0.0$ $18.1\pm 1.5$ $36.4\pm 1.1$ $49.0\pm 3.0$ $88.7\pm 2.3$ 100 $\pm$ 0.0
Alignment $100\pm 0.0$ $31.1\pm 3.1$ $52.0\pm 2.7$ $40.0\pm 2.4$ $96.0\pm 0.8$ 100 $\pm$ 0.0
Parts Assembly $98.9\pm 0.6$ $33.3\pm 1.5$ $53.9\pm 1.1$ $41.3\pm 1.2$ $95.0\pm 0.6$ 100 $\pm$ 0.0
Rearrange $73.3\pm 0.6$ $5.6\pm 1.0$ $7.4\pm 1.1 11$$8\pm 1.8 1.1$$1\pm 0.6$ 64.3 $\pm$ 0.7
Burger Cooking $100\pm 0.0$ $27.8\pm 2.8$ $50.0\pm 4.8$ $48.6\pm 6.9$ $27.8\pm 2.8$ 91.7 $\pm$ 4.8
Bridge Building $100\pm 0.0$ $43.3\pm 3.3$ $53.3\pm 3.8$ $51.7\pm 2.5$ $20.0\pm 0.0$ 87.2 $\pm$ 4.3
Tower of Hanoi $100\pm 0.0$ $14.3\pm 0.0$ $14.3\pm 0.0$ $14.3\pm 0.0$ $14.3\pm 0.0$ 100 $\pm$ 0.0
Overall $95.7\pm 0.1$ $35.7\pm 0.5$ $52.5\pm 0.4$ $48.6\pm 0.8$ $69.9\pm 0.7$ 93.3 $\pm$ 0.7
7 Results
Through the experiments, we aim to answer the following research questions: (1) How does PDDLLM perform relative to other LLM-based planners? (2) Can PDDLLM generalize to unseen, more complex tasks? (3) Does PDDLLM derive high-quality domains with performance comparable to expert designs? (4) How does PDDLLM’s token cost compare to other LLM-based planners? (5) How stable is PDDLLM when faced with suboptimal demonstration?
Q1. Performance comparison to baselines:
Table˜1 presents the planning success rate of all evaluated methods across all tasks, measured with a 50-second time limit. PDDLLM shows clear advantages in planning efficiency and generalizability over baseline methods. While LLM-based baselines perform competitively in simpler tasks like stacking and unstacking, their performance drops sharply in complex tasks such as rearrangement, burger cooking, bridge building, and Tower of Hanoi. In contrast, PDDLLM maintains strong performance across all task categories, achieving an over 40% improvement in overall planning success rate compared to the best LLM-based planner baseline (i.e., LLMTAMP-FF). Even the ablated variant of PDDLLM, RuleAsMem, outperforms the LLM-based planners by at least 17.4% in overall success rate. Compared to PDDLLM, RuleAsMem exhibits less stability. It performs well in simpler tasks but struggles in more complex ones, suggesting the LLMs struggle to understand complex domains. In addition to the results evaluated with a 50-second time window, we also report the overall success rates of the main methods across varying time limits. As shown in Figure˜3(right), PDDLLM consistently outperforms the baseline methods and demonstrates superior planning performance across all time limits. Notably, it reaches performance saturation the fastest, highlighting its superior time efficiency among all evaluated approaches.
We further compare PDDLLM’s planning ability with more powerful reasoning LLMs (OpenAI-o1 and DeepSeek-R1) in Table˜2. While reasoning-based models exhibit strong planning capabilities, their high computational cost prevents them from completing within the 50-second window. To sufficiently compare the planning capabilities, we extend the time limit to 500 seconds and evaluate them on the most challenging tasks. Although o1-TAMP and R1-TAMP show remarkable improvement compared to GPT-4o-based LLMTAMP, they remain less robust in long-horizon planning and fail to match PDDLLM’s performance. In contrast, our method, relying solely on GPT-4o, consistently achieves higher success rates across complex tasks. Detailed time costs for reasoning models are provided in Section˜A.6.
Q2. Generalization:
The experimental results highlight PDDLLM’s ability to handle increasing complexity in both planning and domain derivation, consistently outperforming baseline methods. In terms of planning complexity, as shown in Figure˜3(left), PDDLLM maintains robust planning performance as task complexity grows, achieving high success rates even in scenarios involving up to 20 objects. As shown in Table˜1, performance degradation is observed in more challenging tasks such as rearrangement, where longer action sequences are required and there are more complex motion constraints. From the perspective of domain derivation complexity, PDDLLM remains effective even in tasks demanding the generation of over 100 predicates. However, success rates drop in the most complex domains, such as bridge building, primarily due to missing supporting predicates. More details of the generalization of each task are provided in section˜A.5.
PDDLLM demonstrates a strong ability to integrate knowledge across demonstrations. The modular nature of PDDL action syntax allows the easy transfer of actions learned from different demonstrations among domains. Despite receiving only one example each for stacking and unstacking, it successfully combines the “stack” action and the “unstack” action to solve rearrangement tasks. Similarly, in the bridge building domain, PDDLLM successfully combines the “align” action and the “stack” action using one demonstration of cube alignment and one of stacking. The high success rate of these tasks in Table˜1 underscores the robustness of the derived planning domains.
Figure 3: (left) Planning success rate trend across increasing object counts. (right) Overall planning success rate under varying time limits.
Q3. Domain Quality:
We evaluate the quality of planning domains generated by PDDLLM by comparing the percentage of missing or redundant predicates(actions) and the planning success rate against expert-designed domains. The full definition of missing or redundant can be found in Section˜A.11. None of the generated domains are missing actions. The predicate statistics are reported in Table˜4. The table summarizes the percentage of missing or redundant predicates, omitting tasks without such issues. These results show that PDDLLM produces high-quality domains with minimal errors, even in complex scenarios. While a few predicates may be absent, the overall logical structure remains sound, as reflected in the high overall planning success rate of 93.3% shown in Table˜1, closely matching the performance of expert-crafted domains.
Q4. Token efficiency:
We compare the token efficiency of our method against LLM-based baselines on the three most complex tasks, as shown in Table˜2. Although the GPT-4o-based LLMTAMP incurs lower total token costs, it performs poorly across all three tasks. Compared to o1-TAMP and R1-TAMP, PDDLLM uses significantly fewer tokens while consistently achieving better performance. These results also underscore the challenge of deploying reasoning LLMs on real robot systems, given their high token consumption. Beyond the main research questions, we also evaluated the robustness of our method under variations in prompting styles. Results show that PDDLLM remains stable across different prompts. Full details are provided in Section˜A.7.
Table 2: Comparison of planning success rate (%) and token cost (k) between PDDLLM and LLMTAMP and the reasoning LLM variants. The best results are shown in bold, and the second-best results are underlined.
Task Success Rate (%) $\uparrow$ Token Cost (k) $\downarrow$
PDDLLM LLMTAMP o1-TAMP R1-TAMP PDDLLM LLMTAMP o1-TAMP R1-TAMP
Rearrangement $\mathbf{73.8\pm 1.1}$ $5.6\pm 1.0$ $70.8\pm 1.5$ $40.0\pm 5.0$ 334 212 1200 1460
Tower of Hanoi $\mathbf{100.0\pm 0.0}$ $14.3\pm 0.0$ $33.3\pm 2.4$ $14.3\pm 0.0$ 535 36 529 353
Bridge Building $\mathbf{87.2\pm 4.3}$ $44.3\pm 3.3$ $51.7\pm 2.5$ $40.0\pm 0.0$ 375 50 270 363
Overall $\mathbf{80.5\pm 0.5}$ $13.9\pm 0.9$ $61.5\pm 1.3$ $35.9\pm 3.1$ 415 99 666 725
Q5. Suboptimal demonstrations:
Lastly, we conducted experiments on the bridge-building task to evaluate how PDDLLM adapts to sub-optimal demonstrations by intentionally masking or perturbing input trajectories. The full skill set required for this task consists of pick, stack, and align actions. In the first experiment, only a pick trajectory was provided; subsequent demonstrations incrementally added stack and then align. As shown in Table˜4, success rates improved as demonstrations covered more required actions, reflecting the system’s ability to aggregate knowledge across trajectories. In a fourth experiment, we introduced a noisy trajectory corresponding to an unrelated unstack action. Although this action was initially incorporated into the domain, its negative impact was light: it was never invoked in a solution and was later pruned due to low use frequency.
Table 3: Percentage of missing or redundant predicates and actions across tasks.
Task Missing Redundant
Stack 0.0% 12.5%
Burger Cooking 22.2% 0.0%
Bridge Building 22.2% 0.0%
Tower of Hanoi 0.0% 14.3%
Table 4: Bridge-building success rate (%) under varying demonstration conditions.
Class Demonstration Type Success Rate (%) $\uparrow$
Missing pick $0.0\pm 0.0$
Missing pick & stack $20.0\pm 0.0$
Complete pick & stack & align $86.7\pm 3.8$
Redundant pick & stack & align & unstack $83.3\pm 3.3$
8 Real Robot Deployment
Figure 4: Real-robot experiment in three different platforms
We validate PDDLLM on real robots to demonstrate cross-platform deployability. The system is tested on the Agilex Piper, Franka Panda, and the UR5e, using ArUco markers for pose estimation and ROS2 for control. All platforms successfully complete tasks, including stacking, bridge building, burger cooking, and the Tower of Hanoi through direct deployment of PDDLLM (details in Section˜A.1). We also tested how PDDLLM integrates with learned skills such as VLA: logical actions serve directly as prompts to condition policies, bypassing explicit motion constraints. A real-robot experiment with VLA demonstrates this seamless interface (details in Section˜A.1.3), highlighting the framework’s flexibility across classical and learning-based control.
9 Limitation
A limitation of PDDLLM is its occasional omission of complex predicates, leading to lower success rates in harder domains like Burger Cooking and Bridge Building (Table˜1). For example, in bridge building the surface may be assembled before the base is complete, making subsequent placements infeasible. Expert-designed domains address this by introducing a predicate (all_base_finished) to enforce ordering. Recent work has sought to overcome this limitation by performing domain refining (Liang et al., 2024; Huang et al., 2025). As with many TAMP methods, modeling complex geometric conditions such as deformable objects is challenging in logical space. Nevertheless, we demonstrate that PDDLLM can be integrated with approaches targeting such problems, including VLA methods.
10 Conclusion
This paper presents PDDLLM, the first approach in the field to generate a complete planning domain from scratch, without relying on any predefined predicates or actions. By extracting logical structures directly from pre-trained LLMs, PDDLLM autonomously derives both predicates and actions, enabling fully automated domain construction. Evaluated across a wide range of environments, PDDLLM demonstrates high quality in domain derivation and strong generalizability across diverse task categories. Moreover, when integrated with the LoCA framework, PDDLLM fully automates the integration between the PDDL planning domain and the low-level motion planner. This level of automation significantly improves usability and positions the framework as an adaptable and scalable solution for robotic planning and decision-making. Compared to existing methods, PDDLLM outperforms other LLM-based baselines and closely matches the performance of expert-designed planning domains, particularly in complex and long-horizon planning scenarios. Future work will focus on integrating perception into the system, enabling domain derivation from raw sensory inputs. From the aspect of broader impact, PDDLLM lowers the barrier to deploying robots by enabling domain derivation from demonstrations, making automation more accessible and adaptable.
Appendix A Appendix
A.1 Real Robot Experiment
A.1.1 Franka Panda Arm
Figure 5: Franka Panda Arm building a bridge
Figure 6: Franka Panda Arm solving the Tower of Hannoi puzzle
A.1.2 Agilex Piper Arm
Figure 7: Agilex Piper Arm stacking cubes
Figure 8: Agilex Piper Arm making burgers
A.1.3 VLA Experiment with UR5e
Figure 9: UR5e Arm Stacking Tabletop Objects
Experiment with the vision-language-action model is conducted using a $\pi$0 model  (Black et al., 2024), finetuned on 200 real-world pick and place trajectories. The resulting policy serves as the action module within our pipeline and is invoked by the PDDLLM planning system. In this setup, PDDLLM generates a high-level task plan, while the execution of each logical action in the plan is carried out directly by the Pi0 policy. Importantly, the VLA experiment demonstrates the flexibility of our framework: the action execution layer can be implemented with either a classical motion planner or a learning-based policy, depending on the task requirements.
A.2 Baseline Implementation Details
LLMTAMP: LLM-based task and motion planning (LLMTAMP) builds on methods from (Huang et al., 2022; Li et al., 2022), which use pre-trained LLMs for task planning. We formulate the planning problem as a natural language description. Human demonstrations are interpreted as action sequences required to accomplish the task. The LLM then generates high-level actions to achieve the goal, which are refined into motion plans using predefined skills.
LLMTAMP-FF: LLMTAMP-FF, following the method by (Huang et al., 2023; Chen et al., 2023), extends LLMTAMP with a failure feedback loop. Upon execution failure, the system feeds the failure signal to the LLM to regenerate the task plan, repeating until success or the time limit is reached.
LLMTAMP-FR: Following Wang et al. (2024b), LLMTAMP-FR extends failure detection by providing specific failure reasons to guide replanning with the LLM. We design a reasoner that generates detailed explanations for plan failures and incorporates them into the prompt as feedback. The LLM performs failure reasoning and then regenerates the plan accordingly.
Expert Design: The expert design baseline uses expert-crafted planning domains to evaluate how closely PDDLLM-generated domains approach human-level performance. To highlight the readability and customizability of PDDLLM, these expert-designed domains are initialized with PDDLLM-generated outputs, which are then analyzed and refined by a TAMP expert into ground-truth domains.
o1-TAMP and R1-TAMP: As reasoning models have demonstrated superior performance in many tasks (Zhong et al., 2024), we ablate the LLM in LLMTAMP to compare the planning performance of state-of-the-art reasoning LLMs with our method. Specifically, we evaluate OpenAI’s o1 (OpenAI, 2024) and Deepseek’s R1 (DeepSeek-AI, 2025) as the reasoning backbones.
RuleAsMem: RuleAsMem is an ablation of PDDLLM that treats the generated PDDL domain as contextual memory, rather than using it with a symbolic planner. While prior work focuses on translating language into logical representations (Liu et al., 2023; Xie et al., 2023), RuleAsMem directly integrates logical planning rules into the LLM prompt to solve new tasks. Each task is defined by initial and goal states, using the imagined predicates, along with a human demonstration in the form of a PDDL task plan as prompts.
A.3 Prompt template
A.3.1 Predicate imagination
Template
There are n objects in the environment, whose feature name are feature value 1 and feature value 2. In dimension 1, we know feature subspace range in dimension 1. In dimension 2, we know feature subspace range in dimension 2.Please create a predicate in PDDL syntax to describe this relation and classify if it is related to the current task description. Please return the result in the following format: predicate, relevance.
Note: Here, $u$ is a variable determined by feature subspace range length, which depends on the user input.
A.3.2 Action invention
Template
Current state: The logical state before action execution
Next state:The logical state after the action execution
Can you give this transition an action name to summarize and describe what happened. Please follow Planning Domain Definition Language syntax.
A.3.3 LLMTAMP baseline
Part 1. Initialization:
Imagine you are a robot arm operator; you need to generate a sequence of actions to achieve the given goal. Here are the logical actions you can choose from: list of actions to choose from
Part 2. Human demonstrations:
Here are some examples for you to learn:
example of input and output of the system
Part 3. New planning problem:
Now you are given a new input planning problem as the following: The initial state and the goal of the problem to be solved Choose a sequence of actions to accomplish this task, and return the action sequence following the example output provided.
We integrate parts 1, 2, and 3 into a complete prompt for the LLM to generate task plans.
A.3.4 LLMTAMP+Failure Feedback baseline
LLMTAMP+Failure Feedback extends LLMTAMP with a failure feedback loop. Upon execution failure, the system feeds the failure signal to the LLM for replanning. Thus, the initial prompts of this baseline, part 1, 2, and 3, are the same as the LLMTAMP prompt. However, there is a failure summarization. Integrating part 1-4 gives the full prompt.
Part 4. Failure feedback:
Your plan failed in execution, please generate a different one. Only return the sequence of logical actions following the format of example output.
A.3.5 LLMTAMP+Failure Reasoning baseline
LLMTAMP + Failure Reasoning further extends failure detection by providing specific failure reasons to guide replanning with the LLM. Part 1, 2, and 3 are the same as the LLMTAMP prompt. Integrating part 1-5 gives the full prompt.
Part 4. Failure Reasoning:
Your plan failed in execution, please generate a different one. This may involve sample new plans or reorder the last plan. Please generate output step-by-step, which includes your reasoning for the failure of the last plan. Answer the questions: (i) what is the cause of the failure of the last plan? (ii) do you see similar mistakes in other steps in the plan? Here are the failure reasons: failure reasons
Part 5. Replan:
Now, based on your above failure reasoning for (i) and (ii) , generate a new task plan to solve the original planning problem. Only return the sequence of logical actions following the format of example output.
A.3.6 RuleAsMem
The overall template of RuleAsMem is very similar to that of LLMTAMP. However, in RuleAsMem, the planning domain is provided and the problem is defined in PDDL syntax.
Part 1. Initialization:
Imagine you are a robot arm operator, you need to generate a sequence of actions to achieve the given goal. Here is the PDDL planning domain: PDDL planning domain
Part 2. Human demonstrations:
Here are some examples for you to learn:
example of input and output of the system
Part 3. New planning problem:
Now you are given a new input planning problem as the following: The initial state and the goal of the problem in PDDL syntax. Choose a sequence of actions to accomplish this task. Only return the sequence of logical actions following the format of example output.
A.4 Experiment Tasks
A.4.1 Stacking
The stacking task involves collecting individual objects and placing them on top of each other to form stable stacks.
A.4.2 Unstacking
The unstacking task is the inverse process of stacking, requiring the robot to identify, grasp, and remove items from existing stacks without disturbing surrounding structures.
A.4.3 Rearrangement
The rearrangement task demands the robot to relocate objects from an initial configuration into a desired layout.
A.4.4 Alignment
The alignment task requires the robot to position multiple objects in a straight line with consistent spacing and orientation.
A.4.5 Color classification
In the color classification task, the robot must identify the color of each object, group them by color category, and stack or place them in designated areas accordingly.
A.4.6 Parts assembly
The parts assembly task involves recognizing components and sequentially assembling multiple machining parts together.
A.4.7 Tower of Hanoi
The Tower of Hanoi is a puzzle that involves moving a stack of disks from one base to another, one at a time, without ever placing a larger disk on top of a smaller one, using a third base as an auxiliary.
A.4.8 Bridge building
In bridge building, the robot is required to collect distributed blocks and configure it into a bridge structure.
A.4.9 Burger cooking
Lastly, for burger cooking, the robot needs to stack and pack the food ingredients together to make hamburgers.
A.5 Generalization Across Task Complexity
Figure 10: Planning success rate trend across increasing object counts for each task.
A.6 Time Cost of LLMTAMP Reasoning Model Variants
As noted in the paper, the reasoning models incur substantial computational overhead when generating task plans. To account for this, we extended the time limits for o1-TAMP and R1-TAMP in order to evaluate the contribution of reasoning to planning performance and to enable a fair comparison with our method. This section presents a comprehensive comparison of the time costs between the LLMTAMP variants and our approach.
As shown in Table˜5, our method consistently yields lower average planning times across all tasks compared to the LLMTAMP reasoning model variants. This improvement is primarily attributed to PDDLLM’s ability to structurally summarize the reasoning process into a standardized planning domain during a one-time offline inference step. As a result, no additional reasoning is required at test time. In contrast, LLMTAMP variants conduct reasoning independently for each task instance, leading to significantly higher computational costs. Such overhead makes these models impractical for deployment on physical robots, where real-time planning capabilities are often essential. As shown in Table˜2, while o1-TAMP achieves comparable performance in certain tasks, the substantially higher time cost undermines its practical value, particularly in real robotic scenarios.
Table 5: Comparison of time cost between PDDLLM and LLMTAMP reasoning variants
Experiment Ours o1-TAMP o1-TAMP
Average Planning Time Cost (Second)
Stack 5.94 39.22 211.40
Rearrangement 15.74 92.83 337.10
Tower of Hanoi 4.29 167.82 353.88
Bridge Building 6.45 82.47 305.07
A.7 Prompt Variation Test
In addition to the five main research questions discussed in the paper, we further evaluated the robustness of our method under variations in the prompting styles. Four different cases were tested to assess the stability of domain generation: parallel prompting with varying numbers of prompts, altering the prompting sequence of simulation outcomes, and tuning the prompting template.
A.7.1 Parallel prompting for predicate imagination
In this prompt variation, we perform parallel prompting of the LLM multiple times to obtain diverse responses simultaneously. These outputs are subsequently analyzed and aggregated to synthesize an optimal solution. While the core prompting procedure follows the same structure as described in Section˜A.3.1, additional post-processing steps are applied to summarize and consolidate the results. For prompt naming, we further prompt the LLM to select the most suitable name from among the parallel outputs. The prompt template is shown as the following:
Here is a list of predicate describing the same object state: predicate list. Please choose the PDDL predicate from the provided ones to best describe the scenario. Return the chosen one in PDDL syntax.
For relevance classification, if the parallel outputs are inconsistent, PDDLLM selects the majority response. In the case of a tie, it randomly selects one among the tied options.
A.7.2 Altering the prompting sequence of simulation outcomes
In this prompt variation, we alter the ordering of dimensions presented to the LLM. While the prompt template remains similar to that described in Section˜A.3.1, we do not adhere to a fixed dimension sequence. Instead, we shuffle the order of dimensions within the prompt to evaluate whether this affects the generated output. Some examples are provided here.
A.7.3 Template tuning
In this experiment, we aim to evaluate whether slight modifications to the prompt template affect the LLM’s ability to generate predicate names and assess their relevance. Here is the fine-tuned template:
There are n objects in the environment, whose feature name are feature value 1 and feature value 2. In dimension 1, we know feature subspace range in dimension 1. In dimension 2, we know feature subspace range in dimension 2. Please create a predicate in PDDL syntax to describe this relation. Assign a score to this predicate indicating its relevance to the task of current task description. The score range is 0 to 1, where 0 indicate irrelevant and 1 indicate very relevant. Please return the result in the following format: predicate, score.
After collecting the scores, we set a threshold to determine which predicates are relevant. In the experiment, the threshold chosen is 0.5.
The results, presented in Table˜6, demonstrate that our method remains robust across different prompting styles. Regardless of the prompt variation, the generated planning domains consistently solve the test tasks with a success rate approaching 100%. Minor fluctuations are attributed to randomness in the planning search process and occasional motion execution failures.
Table 6: Planning success rate for domains generated using different prompt styles.
Experiment 10-Parallel 5-Parallel Sequence Altering Template Tuning
Stack 97.8% 96.1% 96.1% 95.6%
Unstack 97.8% 100% 98.9% 98.3%
Cube Alignment 100% 100% 100% 100%
A.8 Planning Time Limit Variation
Table 7: Planning success rate (%) across tasks for all methods (Time limit = 25 s).
Method Expert LLMTAMP LLMTAMP-$F$ LLMTAMP-FR RuleAsMem PDDLLM
Stack $95.8\pm 0.2$ $41.5\pm 4.3$ $56.8\pm 3.4$ $43.8\pm 3.8$ $84.3\pm 3.3$ $97.5\pm 1.6$
Unstack $99.4\pm 0.6$ $81.2\pm 1.1$ $85.0\pm 5.7$ $85.5\pm 5.4$ $79.8\pm 1.9$ $94.9\pm 0.5$
Color Classification $96.3\pm 0.1$ $18.1\pm 1.5$ $24.9\pm 0.8$ $23.1\pm 3.3$ $87.6\pm 1.9$ $99.5\pm 0.4$
Alignment $100.0\pm 0.0$ $31.6\pm 3.1$ $40.9\pm 2.0$ $35.3\pm 2.0$ $96.0\pm 0.8$ $100.0\pm 0.0$
Parts Assembly $98.9\pm 0.6$ $33.6\pm 1.5$ $46.1\pm 2.1$ $37.5\pm 1.4$ $95.1\pm 0.6$ $100.0\pm 0.0$
Rearrange $67.\pm 0.6$ $5.6\pm 1.1$ $11.7\pm 0.6$ $7.4\pm 1.1$ $1.1\pm 0.6$ $52.5\pm 2.2$
Burger Cooking $100.0\pm 0.0$ $27.8\pm 2.8$ $45.1\pm 7.3$ $38.9\pm 5.6$ $27.8\pm 2.8$ $89.6\pm 3.2$
Bridge Building $100.0\pm 0.0$ $43.3\pm 3.3$ $48.9\pm 5.9$ $47.2\pm 4.3$ $20.0\pm 0.0$ $87.2\pm 4.3$
Tower of Hanoi $83.3\pm 2.4$ $14.3\pm 0.0$ $14.3\pm 0.0$ $14.3\pm 0.0$ $14.3\pm 0.0$ $85.7\pm 0.0$
Overall $93.2\pm 0.2$ $34.5\pm 0.4$ $43.2\pm 0.8$ $38.1\pm 0.9$ $68.3\pm 0.9$ $90.5\pm 0.9$
Table 8: Planning success rate (%) across tasks for all methods (Time limit = 100 s).
Method Expert LLMTAMP LLMTAMP-$F$ LLMTAMP-FR RuleAsMem PDDLLM
Stack $99.5\pm 0.5$ $41.5\pm 4.3$ $76.7\pm 2.7$ $71.0\pm 2.7$ $85.5\pm 2.9$ $97.5\pm 1.6$
Unstack $96.1\pm 0.2$ $90.3\pm 1.5$ $96.9\pm 1.2$ $96.1\pm 1.1$ $88.4\pm 1.2$ $97.7\pm 0.7$
Color Classification $100.0\pm 0.0$ $18.1\pm 1.5$ $42.0\pm 1.9$ $64.0\pm 2.2$ $88.7\pm 2.3$ $100.0\pm 0.0$
Alignment $100.0\pm 0.0$ $31.6\pm 3.1$ $55.7\pm 3.8$ $44.3\pm 3.8$ $96.0\pm 0.8$ $100.0\pm 0.0$
Parts Assembly $98.9\pm 0.6$ $33.6\pm 1.5$ $57.1\pm 0.8$ $47.7\pm 2.5$ $95.1\pm 0.6$ $100.0\pm 0.0$
Rearrange $73.9\pm 0.2$ $5.6\pm 1.1$ $17.4\pm 1.1$ $4.7\pm 1.2 1.1$$1\pm 0.6$ $69.4\pm 0.5$
Burger Cooking $100.0\pm 0.0$ $27.8\pm 2.8$ $50.0\pm 4.8$ $51.4\pm 6.1$ $27.8\pm 2.8$ $97.2\pm 2.8$
Bridge Building $100.0\pm 0.0$ $43.3\pm 3.3$ $53.3\pm 3.8$ $57.8\pm 2.2$ $20.0\pm 0.0$ $87.2\pm 4.3$
Tower of Hanoi $100.0\pm 0.0$ $14.3\pm 0.0$ $14.3\pm 0.0$ $14.3\pm 0.0$ $14.3\pm 0.0$ $100.0\pm 0.0$
Overall $95.9\pm 0.1$ $35.8\pm 0.4$ $55.6\pm 1.2$ $54.8\pm 0.7$ $69.9\pm 0.7$ $94.2\pm 0.5$
As outlined in the experimental design section, we set a default planning time limit of 50 seconds to reflect the real-time constraints commonly imposed on robotic systems. Although prior studies typically allow planning times on the order of minutes (Silver et al., 2020; Khodeir et al., 2023; Huang et al., 2025), the exact limits vary considerably. To assess the robustness of our approach under different time allowances, we evaluate performance at time limits of 25, 50, and 100 seconds. The results show that our method consistently outperforms all LLM-based baselines across all tested time settings and task types. Moreover, the planning domains derived by PDDLLM yield performance comparable to that of expert-designed domains under all time settings and across all tasks. All simulations were conducted on a system with an Intel Core i9-14900KF CPU without GPU acceleration, and all LLM prompting was performed using API services.
A.9 Example of PDDLLM imagined predicates
Table 9: Examples of imagined predicates across multiple categories.
Predicate Name
Object Position Predicates
(above ?a ?b) (forall ?a (above ?a ?b))
(forall ?a (above ?a ?b) is false) (not (above ?a ?b))
(beside ?a ?b) (forall ?a (beside ?a ?b))
(forall ?a (beside ?a ?b) is false) (not (beside ?a ?b))
(front ?a ?b) (forall ?a (front ?a ?b))
(forall ?a (front ?a ?b) is false) (not (front ?a ?b))
(far-x ?a ?b) (forall ?a (far-x ?a ?b))
(forall ?a (far-x ?a ?b) is false) (not (far-x ?a ?b))
(apart-y ?a ?b) (forall ?a (apart-y ?a ?b))
(forall ?a (apart-y ?a ?b) is false) (not (apart-y ?a ?b))
(distant-x ?a ?b) (forall ?a (distant-x ?a ?b))
(forall ?a (distant-x ?a ?b) is false) (not (distant-x ?a ?b))
Object Support Predicates
(on-table ?a ?t) (not (on-table ?a ?t))
(forall ?a (on-table ?a ?t)) (forall ?a (on-table ?a ?t) is false)
(aligned-x ?a ?t) (not (aligned-x ?a ?t))
(forall ?a (aligned-x ?a ?t)) (forall ?a (aligned-x ?a ?t) is false)
(aligned-y ?a ?t) (not (aligned-y ?a ?t))
(forall ?a (aligned-y ?a ?t)) (forall ?a (aligned-y ?a ?t) is false)
(above-table ?a ?t) (not (above-table ?a ?t))
(forall ?a (above-table ?a ?t)) (forall ?a (above-table ?a ?t) is false)
(near-x ?a ?t) (not (near-x ?a ?t))
(forall ?a (near-x ?a ?t)) (forall ?a (near-x ?a ?t) is false)
(far-x ?a ?t) (not (far-x ?a ?t))
(forall ?a (far-x ?a ?t)) (forall ?a (far-x ?a ?t) is false)
(far-y ?a ?t) (not (far-y ?a ?t))
(forall ?a (far-y ?a ?t)) (forall ?a (far-y ?a ?t) is false)
(near-each ?a ?t) (not (near-each ?a ?t))
(forall ?a (near-each ?a ?t)) (forall ?a (near-each ?a ?t) is false)
Robot Predicates
(holding ?a ?r) (not (holding ?a ?r))
(forall ?a (holding ?a ?r)) (forall ?a (holding ?a ?r) is false)
(gripper-near-open ?a ?r) (not (gripper-near-open ?a ?r))
(forall ?a (gripper-near-open ?a ?r)) (forall ?a (gripper-near-open ?a ?r) is false)
(gripper-far ?a ?r) (not (gripper-far ?a ?r))
(forall ?a (gripper-far ?a ?r)) (forall ?a (gripper-far ?a ?r) is false)
(gripper-near ?a ?r) (not (gripper-near ?a ?r))
(forall ?a (gripper-near ?a ?r)) (forall ?a (gripper-near ?a ?r) is false)
Color Predicates
similar-color ?a ?b (not similar-color ?a ?b)
(forall ?a similar-color ?a ?b) (forall ?a similar-color ?a ?b is false)
moderate-color-difference ?a ?b (not moderate-color-difference ?a ?b)
(forall ?a moderate-color-difference ?a ?b) (forall ?a moderate-color-difference ?a ?b is false)
distinct-colors ?a ?b (not distinct-colors ?a ?b)
(forall ?a distinct-colors ?a ?b) (forall ?a distinct-colors ?a ?b is false)
very-different-color ?a ?b (not very-different-color ?a ?b)
(forall ?a very-different-color ?a ?b) (forall ?a very-different-color ?a ?b is false)
Size Predicates
(smaller ?a ?b) (not (smaller ?a ?b))
(forall ?a (smaller ?a ?b)) (forall ?a (smaller ?a ?b) is false)
(larger ?a ?b) (not (larger ?a ?b))
(forall ?a (larger ?a ?b)) (forall ?a (larger ?a ?b) is false)
(longer ?a ?b) (not (longer ?a ?b))
(forall ?a (longer ?a ?b)) (forall ?a (longer ?a ?b) is false)
(taller ?a ?b) (not (taller ?a ?b))
(forall ?a (taller ?a ?b)) (forall ?a (taller ?a ?b) is false)
(wider ?a ?b) (not (wider ?a ?b))
(forall ?a (wider ?a ?b)) (forall ?a (wider ?a ?b) is false)
(taller-wider ?a ?b) (not (taller-wider ?a ?b))
(forall ?a (taller-wider ?a ?b)) (forall ?a (taller-wider ?a ?b) is false)
(wider-longer ?a ?b) (not (wider-longer ?a ?b))
(forall ?a (wider-longer ?a ?b)) (forall ?a (wider-longer ?a ?b) is false)
Table 10: Extended examples of imagined predicates for real-world domains.
Predicate Name
Predicates for Weight
Lift a box in warehouse (light-box ?a): 40 < weight < 70 kg
(heavy-box ?a): 70 < weight < 100 kg
Collect apple of different size (small-apple ?a): 100 < w < 166.7 g
(medium-apple ?a): 166.7 < w < 233.3 g
(large-apple ?a): 233.3 < w < 300 g
Predicates for Length
Fulfill a 15-cm coffee cup (needs-refill ?a): 0 < level < 3.75 cm
(half-full ?a): 3.75 < level < 7.5 cm
(nearly-full ?a): 7.5 < level < 11.25 cm
Measuring high school children height (medium-height ?a): 150 < h < 160 cm
(average-height ?a): 160 < h < 170 cm
(tall-height ?a): 170 < h < 180 cm
(very-tall ?a): 180 < h < 190 cm
Predicates for Air Quality
Plan when to travel (suitable-travel ?a): 0 < PM2.5 < 166.7
(moderate-air ?a): 166.7 < PM2.5 < 333.3
(high-risk ?a): 333.3 < PM2.5 < 500
Predicates for Brightness
Fix broken light in office (dimly-lit ?a): 0 < brightness < 150 lux
A.10 Test Problem Set Design
A.10.1 Dimensions of Difficulty
Task Diversity: To ensure robustness and broad applicability, we sampled over 1,200 planning tasks across nine distinct environments, including stacking, unstacking, rearrangement, alignment, parts assembly, color classification, burger cooking, and bridge building. The details of the tasks are provided in Section˜A.4. Each task category includes tasks of 3 to 20 objects, with 10 distinct tasks sampled for each object count. The resulting task plan lengths ranged from 6 to 510 steps, reflecting a wide spectrum of planning horizons. Our experiments spanned over 150 unique predicates.
Task Complexity: In task design, we consider two types of complexity: domain derivation complexity and planning complexity. Domain derivation complexity is determined by the number of predicates imagined and actions invented by PDDLLM; the more predicates and actions, the higher the complexity. Planning complexity is influenced by both the planning domain and the task. Given $n$ objects, a task plan of length $l$, and $m$ actions in the domain, the branching factor at each step is $m\times n$, resulting in a approximate complexity of $(m\times n)^{l}$. In Section˜A.10, we layout a table showcase the planning complexity (in approximate order of magnitude) of the most difficult problem in the category and the domain derivation complexity. The Tower of Hanoi task exhibits the greatest planning complexity, while bridge building presents the highest domain derivation complexity. In contrast, stack and unstack are simpler in both complexity measures.
Knowledge Transferability: We evaluate the knowledge transferability of PDDLLM by testing its ability to generalize from demonstrated tasks to novel ones with overlapping predicates and actions. For most tasks, PDDLLM was given demonstrations of the same task involving fewer (3 to 4) objects. However, for compositional tasks such as rearrangement and bridge building, the model was instead provided demonstrations of simpler subtasks. Specifically, rearrangement used demonstrations from both stacking and unstacking, while bridge building combined demonstrations of stacking and alignment. These setups test whether PDDLLM can compose skills learned from simpler demonstrations to solve more complex tasks.
Table 11: Maximum planning complexity and domain derivation complexity of each category .
Task
Stack
Unstack
Color
Classify
Alignment
Parts
Assembly
Rearrange
Burger
Cook
Bridge
Build
Tower
of Hanoi
Max
Planning
Complexity
$10^{58}$ $10^{58}$ $10^{71}$ $10^{58}$ $10^{58}$ $10^{137}$ $10^{48}$ $10^{36}$ $10^{307}$
Domain
Derivation
Complexity
90 90 111 94 92 96 114 128 94
A.10.2 Derivation of Bounded Domain Derivation Complexity
The predicate imagination complexity measures how many predicates need to be generated. Our method is generalizable to more complex cases for the following reasons:
Complexity calculation formula
The formula is:
$\text{complexity}=\sum_{i}n_{p,i}^{\,n_{\text{dim},i}}$
where $n_{\text{dim},i}$ is the number of dimensions included in predicate type $i$. For example, $n_{\text{dim},\text{color}}=3$ since color is represented by the three RGB channels. $n_{p,i}$ is the number of partitions along each dimension for predicate type $i$. The term $n_{p,i}^{\,n_{\text{dim},i}}$ gives the total number of predicates generated for type $i$, and summing over all predicate types yields the overall complexity of the predicate imagination process for the task.
The Method is Scalable as the Increase in Predicate Types Leads to Linear Growth
The number of predicate types grow linearly as the task complexity increases. The only exponential factor comes from the number of dimensions $n_{\text{dim}}$ per predicate type; however, in practice, this value is typically small (e.g., $\leq 3$).
High-Dimensional Features Can Be Decomposed to Reduce Complexity
For higher-dimensional features, as object pose $(x,y,z,\text{yaw},\text{pitch},\text{roll})$, we can decompose them into lower-dimensional predicate types (e.g., position and orientation). This reduces exponential complexity to additive linear terms, preserving scalability and enabling rich predicate representation.
Task Planning Does Not Require Many Partitions
Dividing features into too many partitions makes predicate implication nearly continuous, which defeats the purpose of symbolic abstraction. Here is an experiment:
  • 2 partitions: “cold” ($0<T<50^{\circ}C$), “hot” ($50<T<100^{\circ}C$)
  • 4 partitions: “cool” ($0<T<25^{\circ}C$), “warm” ($25<T<50^{\circ}C$), “hot” ($50<T<75^{\circ}C$), “near-boiling” ($75<T<100^{\circ}C$)
  • 100 partitions: …, “post-boiling” ($95<T<96^{\circ}C$), “stable-boil” ($96<T<97^{\circ}C$), …
For the same task “boil milk”, we compare the predicate imagination results by forcing the system to partition the feature space into 2, 4, and 100 intervals. It is evident that overly fine partitioning reduces interpretability and semantic efficiency, so the number of partitions $n_{p}$ per dimension is usually low.
A.11 Missing and Redundant Elements
A redundant predicate or action refers to one that, when removed from the planning domain, does not degrade planning performance. Conversely, a missing predicate or action is one whose absence significantly reduces the domain’s planning performance.
The reported percentages are computed by comparing the PDDLLM generated PDDL domains with the expert-designed PDDL domains. The percentage of missing predicates/actions reflects the completeness of the generated domain (i.e., whether all necessary components are present). The percentage of redundant predicates/actions reflects the optimality of the domain (i.e., whether unnecessary components have been excluded).
A.12 Real-world Demonstration Collection
To enable the demonstration collection in real-world, we build up a data collection system using Agilex Pika and ArUco markers. We collect both the end-effector trajectories of the human demonstrator and the pose of the target objects. These data are later mapped to symbolic predicates and actions in the PDDL domain, providing a bridge between raw demonstration data and high-level planning representations.
End-Effector Trajectory: The movement of the human operator’s end effector is recorded using the Agilex Pika Data Collection System. The system’s motion capture capability provides precise position and orientation measurements over time. These trajectories serve as a basis for identifying meaningful segments of motion that can be abstracted into candidate actions for the symbolic domain.
Object Properties and States: Physical characteristics of the objects, such as size, are obtained from the available CAD models, ensuring that the symbolic model is grounded in realistic object dynamics. In addition, the poses of the objects during demonstrations are captured using ArUco markers affixed to their surfaces. These pose estimates allow us to track state changes—such as whether an object has been grasped, moved, or placed—which are subsequently encoded as predicates in the PDDL formulation.